OpenStudy (anonymous):

HELP PLEASE!! Use De Moivre's theorem to express cos5x in terms of powers of sinx and cosx

hartnn (hartnn):

\(\Large \cos 5x +i \sin 5x = (\cos x +i\sin x)^5\) expand right hand side using binomial theorem Real part of right side will be = cos 5x